# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Moving Contact Lines, Variable Contact Angles

### v1

@@ 1   Introduction
In diffuse interface models, the interface between the two materials is represented by a thin layer whose thickness is proportional to a small parameter $\varepsilon>0$. The location of the two components is usually represented by an order parameter, the so-called phase-field, which describes the difference of the local concentrations (or volume fractions) of the materials. This means that the phase-field attains values close to $-1$ or $1$ in the regions where the single fluids are present. At the the diffuse interface between the fluids, the phase-field is expected to exhibit a continuous transition between the values representing the pure phases. In normal direction across the diffuse interface, a certain transition profile can usually be observed. The main advantage of this approach is that the time evolution of the phase-field can be described via an Eulerian formulation as a PDE on the whole domain. This avoids directly tracking the interface between the phases. If a proper scaling with respect to the interface parameter $\varepsilon$ is used, the corresponding sharp-interface model can be recovered (in most cases at least formally) as the sharp-interface limit (see, e.g., [3]).
@@
Based on these considerations, Navier–Stokes–Cahn–Hilliard models with dynamic boundary conditions on the Cahn–Hilliard subsystem as well as a generalized Navier slip boundary condition for the Navier–Stokes equation have been proposed and investigated in the literature. We refer to [40, 41], where a second-order (Allen-Cahn type) dynamic boundary condition was used for the Cahn–Hilliard subsystem, and to [50], where a fourth-order (Cahn–Hilliard type) dynamic boundary condition is considered. The model derived in [50] was further analyzed in [42]. In these models, the Cahn–Hilliard quantities at the boundary enter the generalized Navier slip boundary condition via a forcing term. We also refer to [34, 24], where compressible two-phase flows were considered. In [34] a compressible Navier–Stokes–Allen–Cahn model was analyzed and in [24], a compressible Navier–Stokes–Cahn–Hilliard model with dynamic boundary conditions was investigated. Moreover, models for two-phase flows with phase-transition were studied in [10, 8].
@@ 1.1   Formulation of our model
[…]
@@
$\partial_{t}\big{(}\rho(\phi){\mathbf{u}}\big{)}+\textnormal{div}\,\big{(}{\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big{)}=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0$ in $\quad\text{in }Q_{T},$ (1.1a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu)$ in $\quad\text{in }Q_{T},$ (1.1b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ in $\quad\text{in }Q_{T},$ (1.1c)
$\partial_{t}^{\circ}\psi=\textnormal{div}_{\Gamma}(m_{\Gamma}(\psi)\nabla_{\Gamma}\theta)-\beta m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu$ on $\quad\text{on }S_{T},$ (1.1d)
$\theta=-\delta\Delta_{\Gamma}\psi+\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\partial_{\mathbf{n}}\phi$ on $\quad\text{on }S_{T},$ (1.1e)
$K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi,\quad Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu,\quad K,L\in[0,\infty],$ on $\quad\text{on }S_{T},$ (1.1f)
$\big{[}\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q+\gamma_{\tau}{\mathbf{u}}\big{]}_{\tau}=\Big{[}\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}-\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\Big{]}_{\tau}$ on $\quad\text{on }S_{T},$ (1.1g)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\mathcal{V}-qH$ on $\quad\text{on }S_{T},$ (1.1h)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ on $\quad\text{on }S_{T}.$ (1.1i)
@@
$[\nabla\mathbf{w}]_{ij}=\partial_{j}\mathbf{w}_{i}\quad\text{and}\quad[\textnormal{div}\,\mathbf{M}]_{i}=\sum_{k=1}^{d}\partial_{k}\mathbf{M}_{ik}$
@@
In system (1.1), the volume-averaged velocity field of the two fluids is represented by the vector-valued function ${\mathbf{u}}:Q_{T}\to\mathbb{R}^{d}$. The scalar functions $\phi:Q_{T}\to\mathbb{R}$, $\mu:Q_{T}\to\mathbb{R}$, $\psi:S_{T}\to\mathbb{R}$ and $\theta:S_{T}\to\mathbb{R}$ denote the bulk phase-field, the bulk chemical potential, the surface phase-field, and the surface chemical potential, respectively. It is a key feature of the model (1.1) that the evolving domain $Q_{T}$ is transported by the velocity field ${\mathbf{u}}$. This is reflected in the boundary condition (1.1i)2italic-(1.1isubscriptitalic-)$_{2}$ and, therefore, system (1.1) is a free boundary problem.
@@
$\begin{cases}\phi=\alpha\psi&\text{if $K=0$},\\ K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi&\text{if $K\in(0,\infty)$},\\ \partial_{\mathbf{n}}\phi=0&\text{if $K=\infty$}\end{cases}$ on $\quad\quad\text{on }S_{T},$ (1.6a)
$\begin{cases}\mu=\beta\theta&\text{if $L=0$},\\ Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu&\text{if $L\in(0,\infty)$},\\ m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=0&\text{if $L=\infty$}\end{cases}$ on $\quad\quad\text{on }S_{T}.$ (1.6b)
@@
is used to account for the different choices of the parameter $K$. In (1.7), the first summand on the right-hand side represents the kinetic energy. The second summand is the bulk free energy and the third summand is the surface free energy. They are both of Ginzburg–Landau type and the free energy densities crucially depend on the potentials $F$ and $G$, respectively, which are assumed to be double-well shaped. This means that they are coercive with two local minimum points close to $-1$ and $1$ (as these values represent the pure phases of the materials) and a local maximum point in between (usually at zero). A typical choice for $F$ and $G$ is the so-called Flory–Huggins potential
@@
for $0<\Theta<\Theta_{c}$, which can be derived from Boltzmann kinetics. It is sometimes simply referred to as the logarithmic potential. Since $W_{\textup{log}}^{\prime}(s)\to\pm\infty$ as $s\to\pm\infty$, the potential $W_{\textup{log}}$ is a so-called singular potential. This property ensures that for any solution of system (1.1) with $F=G=W_{\textup{log}}$, the phase-fields $\phi$ and $\psi$ attain their values within the physically relevant range $[-1,1]$. An important consequence is that due to $\left|\phi\right|\leq 1$, we have
@@ 2.1   An alternative reformulation
$\partial_{t}\big{(}\rho(\phi){\mathbf{u}}\big{)}+\textnormal{div}\,\big{(}{\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big{)}-\textnormal{div}\,(\mathbf{S})+\nabla\overline{p}=\mu\nabla\phi,\quad\textnormal{div}\,{\mathbf{u}}=0$ in $\quad\text{in }Q_{T},$ (2.1a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu),$ in $\quad\text{in }Q_{T},$ (2.1b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ in $\quad\text{in }Q_{T},$ (2.1c)
$\partial_{t}^{\circ}\psi=\textnormal{div}_{\Gamma}(m_{\Gamma}(\psi)\nabla_{\Gamma}\theta)-\beta m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu$ on $\quad\text{on }S_{T},$ (2.1d)
$\theta=-\delta\Delta_{\Gamma}\psi+\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\partial_{\mathbf{n}}\phi$ on $\quad\text{on }S_{T},$ (2.1e)
$K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi,\quad Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu,\quad K,L\in[0,\infty],$ on $\quad\text{on }S_{T},$ (2.1f)
$\big{[}\mathbf{S}{\mathbf{n}}+\nabla_{\Gamma}\overline{q}+\gamma_{\tau}{\mathbf{u}}\big{]}_{\tau}=\Big{[}\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}+\theta\nabla_{\Gamma}\psi\Big{]}_{\tau}$ on $\quad\text{on }S_{T},$ (2.1g)
$\mathbf{S}{\mathbf{n}}\cdot{\mathbf{n}}-\overline{p}-\varepsilon(\partial_{\mathbf{n}}\phi)^{2}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\mathcal{V}-\overline{q}H+gH-f$ on $\quad\text{on }S_{T},$ (2.1h)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ on $\quad\text{on }S_{T}.$ (2.1i)
@@ 2.2   Variants of our model without dynamic boundary conditions
$\partial_{t}\big{(}\rho(\phi){\mathbf{u}}\big{)}+\textnormal{div}\,\big{(}{\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big{)}=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0$ in $\quad\text{in }Q_{T},$ (2.2a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu)$ in $\quad\text{in }Q_{T},$ (2.2b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ in $\quad\text{in }Q_{T},$ (2.2c)
$\partial_{\mathbf{n}}\phi=0,\quad\partial_{\mathbf{n}}\mu=0$ on $\quad\text{on }S_{T},$ (2.2d)
$\big{[}\mathbf{S}{\mathbf{n}}+\nabla_{\Gamma}q+\gamma_{\tau}{\mathbf{u}}\big{]}_{\tau}=\Big{[}\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}\Big{]}_{\tau}$ on $\quad\text{on }S_{T},$ (2.2e)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,\mathcal{V}-qH$ on $\quad\text{on }S_{T},$ (2.2f)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ on $\quad\text{on }S_{T},$ (2.2g)
@@
$\partial_{t}\big{(}\rho(\phi){\mathbf{u}}\big{)}+\textnormal{div}\,\big{(}{\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big{)}=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0$ in $\quad\text{in }Q_{T},$ (2.3a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu)$ in $\quad\text{in }Q_{T},$ (2.3b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ in $\quad\text{in }Q_{T},$ (2.3c)
$\partial_{\mathbf{n}}\phi=0,\quad\partial_{\mathbf{n}}\mu=0$ on $\quad\text{on }S_{T},$ (2.3d)
$\big{[}\mathbf{S}{\mathbf{n}}+\gamma_{\tau}{\mathbf{u}}\big{]}_{\tau}=\Big{[}\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}\Big{]}_{\tau}$ on $\quad\text{on }S_{T},$ (2.3e)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,\mathcal{V}$ on $\quad\text{on }S_{T},$ (2.3f)
${\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ on $\quad\text{on }S_{T}.$ (2.3g)

### v2

# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Bulk-Surface Interaction

@@ 1  Introduction
In diffuse interface models, the interface between the two materials is represented by a thin layer whose thickness is proportional to a small parameter $\varepsilon>0$. The location of the two components is usually represented by an order parameter, the so-called phase-field, which describes the difference of the local concentrations (or volume fractions) of the materials. This means that the phase-field attains values close to 1-1 or $1$ in the regions where the single fluids are present. At the the diffuse interface between the fluids, the phase-field is expected to exhibit a continuous transition between the values representing the pure phases. In normal direction across the diffuse interface, a certain transition profile can usually be observed. The main advantage of this approach is that the time evolution of the phase-field can be described via an Eulerian formulation as a PDE on the whole domain. This avoids directly tracking the interface between the phases. If a proper scaling with respect to the interface parameter $\varepsilon$ is used, the corresponding sharp-interface model can be recovered (in most cases at least formally) as the sharp-interface limit (see, e.g., [3]).
@@
Based on these considerations, Navier–Stokes–Cahn–Hilliard models with dynamic boundary conditions on the Cahn–Hilliard subsystem as well as a generalized Navier slip boundary condition for the Navier–Stokes equation have been proposed and investigated in the literature. We refer to [40, 41], where a second-order (Allen-Cahn type) dynamic boundary condition was used for the Cahn–Hilliard subsystem. Here, the Allen-Cahn type dynamic boundary condition can be interpreted as a parabolic relaxation of a transport-type dynamic boundary condition related to contact angle dynamics (see, e.g., [80]). Instead, in [50], a fourth-order (Cahn–Hilliard type) dynamic boundary condition is considered, which allows for a better description of the transfer of material between bulk and boundary. This model was further analyzed in [42]. In these models, the Cahn–Hilliard quantities at the boundary enter the generalized Navier slip boundary condition via a forcing term. We also refer to [34, 24], where compressible two-phase flows were considered. In [34] a compressible Navier–Stokes–Allen–Cahn model was analyzed and in [24], a compressible Navier–Stokes–Cahn–Hilliard model with dynamic boundary conditions was investigated. Moreover, models for two-phase flows with phase-transition were studied in [10, 8].
@@ 1.1  Formulation of our model
${\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}\quad\text{on $S_{T}$}.$
@@
$\partial_{t}\big(\rho(\phi){\mathbf{u}}\big)+\textnormal{div}\,\big({\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big)=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0$ $\quad\text{in }Q_{T},$ (1.1a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu)$ $\quad\text{in }Q_{T},$ (1.1b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ $\quad\text{in }Q_{T},$ (1.1c)
$\partial_{t}^{\circ}\psi=\textnormal{div}_{\Gamma}(m_{\Gamma}(\psi)\nabla_{\Gamma}\theta)-\beta m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu$ $\quad\text{on }S_{T},$ (1.1d)
$\theta=-\delta\Delta_{\Gamma}\psi+\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\partial_{\mathbf{n}}\phi$ $\quad\text{on }S_{T},$ (1.1e)
$K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi,\quad Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu,\quad K,L\in[0,\infty],$ $\quad\text{on }S_{T},$ (1.1f)
$\big[\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q+\gamma_{\tau}{\mathbf{u}}\big]_{\tau}=\Big[\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}-\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\Big]_{\tau}$ $\quad\text{on }S_{T},$ (1.1g)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\mathcal{V}-qH$ $\quad\text{on }S_{T},$ (1.1h)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ $\quad\text{on }S_{T}.$ (1.1i)
@@
$\nabla\mathbf{w}]_{ij}=\partial_{j}\mathbf{w}_{i}\quad\text{and}\quad[\textnormal{div}\,\mathbf{M}]_{i}=\sum_{k=1}^{d}\partial_{k}\mathbf{M}_{ik}$
@@
[…]
@@
$\begin{cases}\phi=\alpha\psi&\text{if $K=0$},\\ K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi&\text{if $K\in(0,\infty)$},\\ \partial_{\mathbf{n}}\phi=0&\text{if $K=\infty$}\end{cases}$ $\quad\quad\text{on }S_{T},$ (1.6a)
$\begin{cases}\mu=\beta\theta&\text{if $L=0$},\\ Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu&\text{if $L\in(0,\infty)$},\\ m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=0&\text{if $L=\infty$}\end{cases}$ $\quad\quad\text{on }S_{T}.$ (1.6b)
@@
is used to account for the different choices of the parameter $K$. In (1.7), the first summand on the right-hand side represents the kinetic energy. The second summand is the bulk free energy and the third summand is the surface free energy. They are both of Ginzburg–Landau type and the free energy densities crucially depend on the potentials $F$ and $G$, respectively, which are assumed to be double-well shaped. This means that they are coercive with two local minimum points close to 1-1 and $1$ (as these values represent the pure phases of the materials) and a local maximum point in between (usually at zero). A typical choice for $F$ and $G$ is the so-called Flory–Huggins potential
@@
for $0<\Theta<\Theta_{c}$, which can be derived from Boltzmann kinetics. It is sometimes simply referred to as the logarithmic potential. Since $W_{\textup{log}}^{\prime}(s)\to\pm\infty$ as $s\to\pm 1$, the potential $W_{\textup{log}}$ is a so-called singular potential. This property ensures that for any solution of system (1.1) with $F=G=W_{\textup{log}}$, the phase-fields $\phi$ and $\psi$ attain their values within the physically relevant range [1,1][-1,1]. An important consequence is that due to $\left|\phi\right|\leq 1$, we have
@@ 2.1  An alternative reformulation
$\partial_{t}\big(\rho(\phi){\mathbf{u}}\big)+\textnormal{div}\,\big({\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big)-\textnormal{div}\,(\mathbf{S})+\nabla\overline{p}=\mu\nabla\phi,\quad\textnormal{div}\,{\mathbf{u}}=0$ $\quad\text{in }Q_{T},$ (2.1a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu),$ $\quad\text{in }Q_{T},$ (2.1b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ $\quad\text{in }Q_{T},$ (2.1c)
$\partial_{t}^{\circ}\psi=\textnormal{div}_{\Gamma}(m_{\Gamma}(\psi)\nabla_{\Gamma}\theta)-\beta m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu$ $\quad\text{on }S_{T},$ (2.1d)
$\theta=-\delta\Delta_{\Gamma}\psi+\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\partial_{\mathbf{n}}\phi$ $\quad\text{on }S_{T},$ (2.1e)
$K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi,\quad Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu,\quad K,L\in[0,\infty],$ $\quad\text{on }S_{T},$ (2.1f)
$\big[\mathbf{S}{\mathbf{n}}+\nabla_{\Gamma}\overline{q}+\gamma_{\tau}{\mathbf{u}}\big]_{\tau}=\Big[\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}+\theta\nabla_{\Gamma}\psi\Big]_{\tau}$ $\quad\text{on }S_{T},$ (2.1g)
$\mathbf{S}{\mathbf{n}}\cdot{\mathbf{n}}-\overline{p}-\varepsilon(\partial_{\mathbf{n}}\phi)^{2}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\mathcal{V}-\overline{q}H+gH-f$ $\quad\text{on }S_{T},$ (2.1h)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ $\quad\text{on }S_{T}.$ (2.1i)
@@ 2.2  Variants of our model without dynamic boundary conditions
$\partial_{t}\big(\rho(\phi){\mathbf{u}}\big)+\textnormal{div}\,\big({\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big)=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0$ $\quad\text{in }Q_{T},$ (2.2a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu)$ $\quad\text{in }Q_{T},$ (2.2b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ $\quad\text{in }Q_{T},$ (2.2c)
$\partial_{\mathbf{n}}\phi=0,\quad\partial_{\mathbf{n}}\mu=0$ $\quad\text{on }S_{T},$ (2.2d)
$\big[\mathbf{S}{\mathbf{n}}+\nabla_{\Gamma}q+\gamma_{\tau}{\mathbf{u}}\big]_{\tau}=\Big[\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}\Big]_{\tau}$ $\quad\text{on }S_{T},$ (2.2e)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,\mathcal{V}-qH$ $\quad\text{on }S_{T},$ (2.2f)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ $\quad\text{on }S_{T},$ (2.2g)
@@
$\partial_{t}\big(\rho(\phi){\mathbf{u}}\big)+\textnormal{div}\,\big({\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big)=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0$ $\quad\text{in }Q_{T},$ (2.3a)
$\partial_{t}^{\bullet}\phi=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu)$ $\quad\text{in }Q_{T},$ (2.3b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi)$ $\quad\text{in }Q_{T},$ (2.3c)
$\partial_{\mathbf{n}}\phi=0,\quad\partial_{\mathbf{n}}\mu=0$ $\quad\text{on }S_{T},$ (2.3d)
$\big[\mathbf{S}{\mathbf{n}}+\gamma_{\tau}{\mathbf{u}}\big]_{\tau}=\Big[\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}\Big]_{\tau}$ $\quad\text{on }S_{T},$ (2.3e)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}+\gamma_{\mathbf{n}}\mathcal{V}=\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,\mathcal{V}$ $\quad\text{on }S_{T},$ (2.3f)
${\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}$ $\quad\text{on }S_{T}.$ (2.3g)